OpenStudy (anonymous):

If you travel 199.46 miles in 1 hour. How many miles will you travel in 2.5 hours?

OpenStudy (anonymous):

199.46 x 2.5

OpenStudy (anonymous):

Okay so its 498.65 thats what i thought

OpenStudy (anonymous):

Thank you so much! :D

OpenStudy (tylerd):

\[\frac{ 199.46miles }{ hour } \times \frac{ 2.5 hours}{ 1 }\]

OpenStudy (tylerd):

hours cancel out and your left with miles